A body of mass m falls from rest through a medium which exerts a frictional drag force bv proportional to a frictional force F. Expand the velocity in power series up to t^4, at what time does the body reach 1/e of it terminal velocity, find it displacement, expand the displacement in power series up to t^4


1
Expert's answer
2021-06-03T09:17:44-0400

"F=mg-bv,"

"m\frac{dv}{dt}=mg-bv,"

"\frac{dv}{dt}=\frac{mg-bv}{m},"

"\frac{dv}{mg-bv}=\frac{dt}{m},"

"v=\frac{mg}{b}(1-e^{-\frac bm t})."